### v2

# No time to train! Training-Free Reference-Based Instance Segmentation

Abstract
The performance of image segmentation models has historically been constrained by the high cost of collecting large-scale annotated data. The Segment Anything Model (SAM) alleviates this original problem through a promptable, semantics-agnostic, segmentation paradigm and yet still requires manual visual-prompts or complex domain-dependent prompt-generation rules to process a new image. Towards reducing this new burden, our work investigates the task of object segmentation when provided with, alternatively, only a small set of reference images. Our key insight is to leverage strong semantic priors, as learned by foundation models, to identify corresponding regions between a reference and a target image. We find that correspondences enable automatic generation of instance-level segmentation masks for downstream tasks and instantiate our ideas via a multi-stage, training-free method incorporating (1) memory bank construction; (2) representation aggregation and (3) semantic-aware feature matching. Our experiments show significant improvements on segmentation metrics, leading to state-of-the-art performance on COCO FSOD (36.8% nAP), PASCAL VOC Few-Shot (71.2% nAP50) and outperforming existing training-free approaches on the Cross-Domain FSOD benchmark (22.4% nAP).
Website: https://miquel-espinosa.github.io/no-time-to-train
1 Introduction
Figure 1: Cross-domain 1-shot segmentation results using our training-free method on CD-FSOD benchmark. Our method directly evaluates on diverse datasets without any fine-tuning, using frozen SAMv2 and DINOv2 models. The reference set contains a single example image per class. The model then segments the entire target dataset based on the reference set. Results show: (1) generalization capabilities to out-of-distribution domains (e.g., underwater images, cartoons, microscopic textures); (2) state-of-the-art performance in 1-shot segmentation without training or domain adaptation; (3) limitations in cases with ambiguous annotations or highly similar classes (e.g., “harbor” vs. “ships” in DIOR). Best viewed when zoomed in. Ablation studies further investigate the variance associated with the selection of reference images. Refer to the Supplementary Material for more visualisations per dataset.
It is well understood that collecting large-scale annotations for segmentation tasks is a costly and time-consuming process [2, 57, 17]. Recent advances in promptable segmentation frameworks [89, 73, 29, 74, 56, 44], epitomised by the Segment Anything Model (SAM) [33, 61], have significantly reduced manual effort by enabling high-quality mask generation using simple geometric prompts such as points, boxes or rough sketches. While this represents a substantial advancement in reducing manual effort, these masks lack semantic awareness [64, 9, 21, 27] and require either manual intervention or complex, domain-specific prompt-generation pipelines to function autonomously (e.g., medical imaging [48, 37, 84, 83], agriculture [5, 69], remote sensing [55, 49, 70]). Relying on manual prompts for each image limits their scalability (especially for large datasets or scenarios requiring automatic processing), while relying on domain-constrained automated pipelines restricts the ability to generalise to cross-domain scenarios.
Reference-based instance segmentation [53, 15, 19] offers a promising solution to this challenge by using a small set of annotated reference images to guide the segmentation of a large set of target images. This idea has the potential to enable cheap, quick and automatic annotation of datasets where such labelling is expensive, time-consuming and requires expertise knowledge [2, 57]. Unlike slow manual prompting [24], using reference images can incorporate semantic understanding directly from examples, and is thus well-suited for automated segmentation tasks. Despite promising results, we observe that existing reference-based segmentation methods often require fine-tuning on novel classes and this raises a set of well understood concerns that include task-specific data requirements, overfitting and domain shift. We conjecture that a prospective alternative approach to guiding reference-based instance segmentation involves reusing the general purpose capabilities of vision foundation models [4, 60, 54, 33, 61].
Several works [46, 67, 74] have attempted to combine pretrained models for reference-based segmentation, e.g. Matcher [46] which integrates DINOv2 with SAM for semantic segmentation tasks. However, these methods face several limitations. Firstly, they rely on computationally expensive distance metrics (Earth Mover’s Distance), and complex thresholding mechanisms, which significantly slow down inference. Secondly, they are not suited for instance-level segmentation tasks, struggling with fine-grained discrimination in complex multi-object scenes. In fact, the instance segmentation setting presents unique challenges—how do we handle occlusions, scale variations, ambiguous object boundaries and varying image quality, all with just a few reference images—and they should be tackled thoughtfully. Effectively combining foundation models, without significant finetuning, remains a large challenge [9], particularly when attempting to leverage generalisation capabilities of semantic ViT backbones (e.g. DINOv2), to achieve precise localisation [82].
We propose a training-free three-stage method: (1) constructing a memory bank of category-specific features, (2) refining feature representations via two-step aggregation, and (3) performing inference through feature matching and a novel semantic-aware soft merging strategy.
This results in a training-free, high-performing framework that achieves significant performance gains on established datasets. Furthermore, our approach maintains its effectiveness across diverse domains with fixed hyperparameters, making it accessible for a wide range of applications. Our contributions can be summarised as:
  • We propose a training-free method that effectively integrates semantic-agnostic segmentation mask proposals with fine-grained semantics for reference-based instance segmentation.
  • We introduce a novel three-stage framework for instance segmentation with vision foundation models, addressing key integration challenges through (1) memory bank construction, (2) two-step feature aggregation, and (3) feature matching with semantic-aware soft merging.
  • Our method achieves state-of-the-art performance on COCO-FSOD, PASCAL-FSOD, and CD-FSOD benchmarks, demonstrating strong generalisation across diverse datasets under fixed hyperparameter settings, without the need for intermediate fine-tuning.
2 Related Work
Figure 2: Overview of our training-free method for few-shot instance segmentation and object detection. (1) Reference Memory Creation: A segmented reference image is processed using the DINOv2 model to generate semantic feature embeddings. (2) Feature aggregation: We compute instance-wise feature representations, and then, aggregate them into class-wise prototypes, stored in the memory bank. (3) Inference on Target Dataset: For each target image, SAMv2 generates instance segmentation masks while DINOv2 extracts semantic features. Using cosine similarity, each mask’s embedding is compared with the reference memory bank to assign the most similar class label. Finally, predictions are aggregated via semantic-aware soft merging to produce the final annotated image. This pipeline enables semantic prompting via reference images, without requiring fine-tuning, and demonstrates state-of-the-art performance on established benchmarks (COCO-FSOD, PASCAL-FSOD) and strong generalization across domains (CD-FSOD).
Reference-based Instance Segmentation aims to segment individual objects within an image, distinguishing between instances of the same category [6, 39]. Traditional approaches, such as Mask R-CNN [22], use region proposals with convolutional networks to predict instance masks, while transformer-based models like DETR [3] and Mask2Former [7] integrate global context through self-attention mechanisms. These methods have demonstrated success in standard instance segmentation tasks by leveraging large, labeled datasets [42, 17]. Reference-based instance segmentation extends these tasks to handle novel categories with limited labeled examples. Early works [15, 53] adapted Mask R-CNN by introducing instance-level discriminative features [15] or uncertainty-guided bounding box prediction [53]. More recent works have unified segmentation tasks into in-context learning frameworks [73, 74, 29], involving an expensive pretraining phase on a wide range of segmentation tasks, including instance segmentation [73, 74] and using contrastive pretraining to integrate visual and textual prompts [29]. Despite these advancements, reference-based instance segmentation remains challenging due to the lack of labeled data, the complexity of multi-instance scenarios, limited generalisation across domains for specialist models, reference image ambiguity, and reliance on predefined class labels [79]. Additionally, reusing frozen backbones not originally pretrained for instance segmentation remains a challenge [82, 9]. Our method effectively reuses two existing frozen vision foundation models—none of which was trained for reference-based instance segmentation task— to tackle reference-based instance segmentation without additional training, while also generalising well to unusual domains.
Vision Foundation models have revolutionised computer vision by learning strong, pretrained representations, transferable across diverse tasks. CLIP [60] and DINO models [4, 54] exemplify this trend, using contrastive learning to align visual and textual representations, and learning robust image embeddings from unlabeled data, respectively. These models have been widely adopted for downstream tasks, including open-vocabulary detection [64, 21] and semantic segmentation [62, 88]. However, CLIP struggles with detailed spatial reasoning, while DINOv2, despite capturing fine-grained semantics, produces low-resolution feature maps. The Segment Anything Models (SAM, SAM2) [33, 61] are a notable addition to this category, trained on an extensive, category-agnostic dataset (SA-1B) [33]. While SAM excels in generating segmentation masks with minimal input (e.g., points, bounding boxes), it lacks inherent semantic understanding [64, 9, 21, 27]. Efforts to bridge this gap include pairing SAM with language models [34], diffusion models [87], or fine-tuning it on labeled datasets like COCO [42] and ADE20K [38]. However, these adaptations often result in complex pipelines or limited scalability. Furthermore, SAM’s semantic-agnostic nature poses limitations in scenarios requiring class differentiation at instance-level. Our work combines the complementary strengths of DINOv2 and SAM without requiring finetuning. By aggregating and matching features from multiple references, we enable high-precision, training-free instance segmentation, achieving state-of-the-art results on diverse few-shot benchmarks.
Automatic Vision Prompting for SAM aims to construct automatic prompting pipelines to enhance SAM’s versatility in complex visual tasks, reducing its reliance on manual inputs. Training-free methods [81, 46] leverage feature-matching techniques but often rely on manually tuned thresholds, distance metrics and complex pipelines. Other approaches focus on learning prompts directly, such as spatial or semantic optimisation [28, 67, 26], but these methods still face challenges in multi-instance or semantically dense settings. Zero-shot methods [78, 65] introduce visual markers to guide attention during segmentation, but lack fine-grained precision. Recent efforts, including SEEM [89], have unified segmentation and recognition tasks through shared decoders, while SINE [45] tackles task ambiguity by disentangling segmentation tasks. Others have paired SAM with Stable Diffusion for open-vocabulary segmentation [87], enabling it to incorporate semantic cues. Similarly, LISA [34] utilises language-based instructions to adapt SAM for text-guided tasks. Despite these efforts, handling multi-instance scenarios and semantic ambiguity without training remains challenging. Our method differs by integrating SAM with DINOv2 without additional training or prompt optimisation. We introduce a multi-stage method (memory bank construction, representation aggregation, and semantic-aware feature matching) with fixed hyperparameters across all experimental settings, making it practical for a wide range of downstream applications and directly accessible to practitioners across diverse domains.
3 Method
3.1 Preliminaries
Segment Anything Model (SAM) [33] is designed for promptable segmentation, that is, it responds to different types of geometric prompts to generate image segmentation masks. It consists of three main components: an image encoder, a prompt encoder, and a mask decoder. The image encoder is a pretrained Vision Transformer (ViT) [8] adapted for high-resolution inputs [40]. The prompt encoder handles both sparse (points, boxes, text) and dense (rough mask) prompts, which are encoded with positional encodings [68], learnt embeddings, and off-the-shelf text encoders. The mask decoder efficiently generates masks by using a modified Transformer decoder block [3] with self-attention and cross-attention on the prompts. For training, SAM uses a combination of focal loss [43] and dice loss [50].
DINOv2 Self-supervised pretrained vision encoder [54] is a self-supervised vision model designed to produce general-purpose visual features. It uses a discriminative self-supervised learning approach based on Vision Transformer (ViT) architectures [8]. DINOv2 makes use of teacher-student networks, incorporates Sinkhorn-Knopp normalisation, multi-crops strategy, separate projection heads for image-level and patch-level objectives, and additional regularisation techniques to stabilise and scale training [4, 85]. It is trained on a curated dataset (LVD-142M) of 142 million images using efficient training techniques (fast memory-efficient attention, stochastic depth, etc.). Its features are transferable across tasks and domains, making it a robust backbone for both global and local visual understanding tasks.
Reference-based instance segmentation aims to segment a target image by using a reference segmented image as an example (rather than using geometric prompts). More formally, given a reference image $I_{r}$ and its corresponding annotations $M^{i}_{r}$ (where $i$ denotes the different object categories), we use this data to segment the corresponding regions in the target image $I_{t}$ that also belong to category $i$.
3.2 Training-free method
The goal of our training-free method is to extract category-specific features from a set of annotated reference examples and use them to segment and classify instances in target images. Unlike methods that require model retraining, we employ a memory-based approach to store discriminative representations for object categories. It consists of three main stages: (1) constructing a memory bank from reference images, (2) refining these representations via two-stage feature aggregation, and (3) performing inference on the target images through feature matching and semantic-aware soft merging. Figure 2 illustrates the complete pipeline.
(1) Memory Bank Construction.
Given a set of reference images $\{I_{r}^{j}\}_{j=1}^{N_{r}}$ and their corresponding instance masks $\{M_{r}^{j,i}\}_{j=1}^{N_{r}}$ for category $i$, we extract dense feature maps $F_{r}^{j}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times d}$ using a pretrained frozen encoder $\mathcal{E}$ [^1] https://github.com/facebookresearch/dinov2, where $d$ is the feature dimension and $H^{\prime},W^{\prime}$ denote the spatial resolution of the feature map. The corresponding instance masks $M_{r}^{j,i}\in\{0,1\}^{H^{\prime}\times W^{\prime}}$ are resized to match this resolution.
For each category $i$, we store the masked features:
$\mathcal{F}_{r}^{j,i}=F_{r}^{j}\odot M_{r}^{j,i}$
where $\odot$ denotes element-wise multiplication. These category-wise feature sets are stored in a memory bank $\mathcal{M}_{i}$, which is synchronised across GPUs to ensure consistency in distributed settings.
(2) Two-stage feature aggregation.
To construct category prototypes, we first compute instance-wise feature representations, and then, aggregate them into class-wise prototypes.
  1. (a)
    Instance-wise prototypes:
    Each instance $k$ in reference image $I_{r}^{j}$ has its own prototype, computed by averaging the feature embeddings within its corresponding mask:
    $P_{r}^{j,k}=\frac{1}{\|M_{r}^{j,k}\|_{1}}\sum_{(u,v)}M_{r}^{j,k}(u,v)F_{r}^{j}(u,v)$
    where $P_{r}^{j,k}\in\mathbb{R}^{d}$ represents the mean feature representation of the $k$-th instance in image $I_{r}^{j}$.
  2. (b)
    Class-wise prototype:
    We compute the category prototype $P_{i}$ by averaging all instance-wise prototypes belonging to the same category $i$:
    $P_{i}=\frac{1}{N_{i}}\sum_{j=1}^{N_{r}}\sum_{k\in\mathcal{K}_{i}^{j}}P_{r}^{j,k},$
    where $\mathcal{K}_{i}^{j}$ is the set of instances in image $I_{r}^{j}$ that belong to category $i$, and $N_{i}=\sum_{j=1}^{N_{r}}|\mathcal{K}_{i}^{j}|$ is the total number of instances belonging to category $i$. These class-wise prototypes $P_{i}$ are stored in the memory bank.
(3) Inference on Target Images.
For a target image $I_{t}$, we extract dense features $F_{t}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times d}$ using the same encoder $\mathcal{E}$. We use the frozen SAM model to generate $N_{m}$ candidate instance masks $\{M_{t}^{m}\}_{m=1}^{N_{m}}$, where $M_{t}^{m}\in\{0,1\}^{H^{\prime}\times W^{\prime}}$. Each mask $M_{t}^{m}$ is used to compute a feature representation via average pooling and L2 normalisation:
$P_{t}^{m}=\frac{1}{\|M_{t}^{m}\|_{1}}\sum_{(u,v)}M_{t}^{m}(u,v)F_{t}(u,v),\quad\hat{P}_{t}^{m}=\frac{P_{t}^{m}}{\|P_{t}^{m}\|_{2}}$
where $\hat{P}_{t}^{m}\in\mathbb{R}^{d}$ is the normalised mask feature.
To classify each candidate mask, we compute:
  1. (a)
    Feature Matching. We compute the cosine similarity between $\hat{P}_{t}^{m}$ and category prototypes $P_{i}$:
    $S_{t}^{m}=\max_{i}\left(\frac{\hat{P}_{t}^{m}\cdot P_{i}}{\|P_{i}\|_{2}}\right)$
    which provides the classification score $S_{t}^{m}$ for mask $M_{t}^{m}$.
  2. (b)
    Semantic-Aware Soft Merging. To handle overlapping predictions, we introduce a novel soft merging strategy. Given two masks $M_{t}^{m}$ and $M_{t}^{m^{\prime}}$ of the same category, we compute their intersection-over-self (IoS):
    $\text{IoS}(M_{t}^{m},M_{t}^{m^{\prime}})=\frac{\sum(M_{t}^{m}\cap M_{t}^{m^{\prime}})}{\sum M_{t}^{m}}$
    and weight it by feature similarity:
    $w_{m,m^{\prime}}=\frac{\hat{P}_{t}^{m}\cdot\hat{P}_{t}^{m^{\prime}}}{\|\hat{P}_{t}^{m^{\prime}}\|_{2}}$
    The final score for each mask is adjusted using a decay factor:
    $S_{t}^{m}\leftarrow S_{t}^{m}\cdot\sqrt{(1-\text{IoS}(M_{t}^{m},M_{t}^{m^{\prime}})w_{m,m^{\prime}}})$
    reducing redundant detections while preserving distinct instances that may partially overlap. Finally, we rank masks by their adjusted scores, and select the top-$K$ predictions as the final output.
3.3 Technical implementation details
Our codebase builds upon SAM2-L (Hierarchical ViT) for mask generation and DINOv2-L as the feature encoder. The encoder processes images at 518×518 resolution with a patch size of 14×14, while SAM2 operates at 1024×1024 resolution. During inference, SAM2 first generates candidate masks using a 32×32 grid of query points. For each mask, we compute L2-normalised features by average pooling the encoder features within the masked region. These features are compared against our memory bank, which stores features from $n$ reference images per category. We employ non-maximum suppression with an IoU threshold of 0.5, followed by our semantic-aware soft merging strategy to handle overlapping predictions. The model outputs up to 100 instances per image. The implementation uses PyTorch [1] and PyTorch Lightning [11] for distributed workload across GPUs.
4 Results
4.1 Object Detection and Instance Segmentation
Although our training-free method outputs segmentation masks, we convert instance masks to bounding boxes for a fair comparison with existing methods.
COCO-FSOD Benchmark
We evaluate our method in a strict few-shot setting on the COCO-20i dataset [42, 30], using the standard 10-shot [^2] https://paperswithcode.com/sota/few-shot-object-detection-on-ms-coco-10-shot and 30-shot [^3] https://paperswithcode.com/sota/few-shot-object-detection-on-ms-coco-30-shot settings. Results for the COCO-FSOD benchmark are shown in Table 1. All results are reported for COCO-NOVEL classes. Novel classes are the COCO categories that intersect with PASCAL VOC categories [10].
Our method achieves state-of-the-art while being completely training-free, outperforming approaches that fine-tune on novel classes.
Figure 3 presents qualitative results, showing our method’s ability to handle multiple overlapping instances in crowded scenes with fine-grained semantics and precise localisation. With semantic-aware soft merging, we mitigate duplicate detections and false positives. Failure cases are discussed in the Supplementary Material.
Method Ft. on novel 10-shot 30-shot nAP nAP50 nAP75 nAP nAP50 nAP75 TFA [72] 10.0 19.2 9.2 13.5 24.9 13.2 FSCE [66] 11.9 10.5 16.4 16.2 Retentive RCNN [13] 10.5 19.5 9.3 13.8 22.9 13.8 HeteroGraph [18] 11.6 23.9 9.8 16.5 31.9 15.5 Meta F. R-CNN [19] 12.7 25.7 10.8 16.6 31.8 15.8 LVC [32] 19.0 34.1 19.0 26.8 45.8 27.5 C. Transformer [20] 17.1 30.2 17.0 21.4 35.5 22.1 NIFF [16] 18.8 20.9 DiGeo [47] 10.3 18.7 9.9 14.2 26.2 14.8 CD-ViTO (ViT-L) [14] 35.3 54.9 37.2 35.9 54.5 38.0 FSRW [31] 5.6 12.3 4.6 9.1 19.0 7.6 Meta R-CNN [77] 6.1 19.1 6.6 9.9 25.3 10.8 DE-ViT (ViT-L) [82] 34.0 53.0 37.0 34.0 52.9 37.2 Training-free (ours) 36.6 54.1 38.3 36.8 54.5 38.7
Table 1: Comparison of our training-free method against state-of-the-art approaches on the COCO-FSOD benchmark under 10-shot and 30-shot settings. Our approach achieves state-of-the-art performance without finetuning on novel classes (Ft. on novel). Results are reported in terms of nAP, nAP50, and nAP75. nAP refers to mAP for novel classes. Competing methods results are sourced from [14]. Since we are the only method that provides both bounding box and segmentation results, for simplicity we omit segmentation AP on this table.
Method Ft. on novel Novel Split 1 Novel Split 2 Novel Split 3 Avg
1 2 3 5 10 1 2 3 5 10 1 2 3 5 10
FsDetView [75] 25.4 20.4 37.4 36.1 42.3 22.9 21.7 22.6 25.6 29.2 32.4 19.0 29.8 33.2 39.8 29.2
TFA [72] 39.8 36.1 44.7 55.7 56.0 23.5 26.9 34.1 35.1 39.1 30.8 34.8 42.8 49.5 49.8 39.9
Retentive RCNN [13] 42.4 45.8 45.9 53.7 56.1 21.7 27.8 35.2 37.0 40.3 30.2 37.6 43.0 49.7 50.1 41.1
DiGeo [47] 37.9 39.4 48.5 58.6 61.5 26.6 28.9 41.9 42.1 49.1 30.4 40.1 46.9 52.7 54.7 44.0
HeteroGraph [18] 42.4 51.9 55.7 62.6 63.4 25.9 37.8 46.6 48.9 51.1 35.2 42.9 47.8 54.8 53.5 48.0
Meta Faster R-CNN [19] 43.0 54.5 60.6 66.1 65.4 27.7 35.5 46.1 47.8 51.4 40.6 46.4 53.4 59.9 58.6 50.5
CrossTransformer [20] 49.9 57.1 57.9 63.2 67.1 27.6 34.5 43.7 49.2 51.2 39.5 54.7 52.3 57.0 58.7 50.9
LVC [32] 54.5 53.2 58.8 63.2 65.7 32.8 29.2 50.7 49.8 50.6 48.4 52.7 55.0 59.6 59.6 52.3
NIFF [16] (*) 62.8 67.2 68.0 70.3 68.8 38.4 42.9 54.0 56.4 54.0 56.4 62.1 61.2 64.1 63.9 59.4
Multi-Relation Det [12] 37.8 43.6 51.6 56.5 58.6 22.5 30.6 40.7 43.1 47.6 31.0 37.9 43.7 51.3 49.8 43.1
DE-ViT (ViT-S/14) [82] 47.5 64.5 57.0 68.5 67.3 43.1 34.1 49.7 56.7 60.8 52.5 62.1 60.7 61.4 64.5 56.7
DE-ViT (ViT-B/14) [82] 56.9 61.8 68.0 73.9 72.8 45.3 47.3 58.2 59.8 60.6 58.6 62.3 62.7 64.6 67.8 61.4
DE-ViT (ViT-L/14) [82] 55.4 56.1 68.1 70.9 71.9 43.0 39.3 58.1 61.6 63.1 58.2 64.0 61.3 64.2 67.3 60.2
Training-free (ours) 70.8 72.3 73.3 77.2 79.1 54.5 67.0 76.3 75.9 78.2 61.1 67.9 71.3 70.8 72.6 71.2
Table 2: AP50 results on the novel classes of the Pascal VOC few-shot benchmark. Competing method results are sourced from [82]. State-of-the-art results are highlighted in bold. (*) indicates that the corresponding implementation is not publicly accessible. Our proposed training-free approach consistently achieves superior performance across all splits, outperforming fine-tuned methods.
[…]
Methods Ft. on novel 1-shot 5-shot $\mathbf{20^{0}}$ $\mathbf{20^{1}}$ $\mathbf{20^{2}}$ $\mathbf{20^{3}}$ mean $\mathbf{20^{0}}$ $\mathbf{20^{1}}$ $\mathbf{20^{2}}$ $\mathbf{20^{3}}$ mean DiffewS [87] 47.7 56.4 51.9 48.7 51.2 52.0 63.0 54.5 54.3 56.0 DiffewS-n [87] 47.1 56.6 53.8 48.3 52.2 57.3 66.5 60.3 58.8 60.7 HSNet [51] 37.2 44.1 42.4 41.3 41.2 45.9 53.0 51.8 47.1 49.5 CyCTR [80] 38.9 43.0 39.6 39.8 40.3 41.1 48.9 45.2 47.0 45.6 VAT [23] 39.0 43.8 42.6 39.7 41.3 44.1 51.1 50.2 46.1 47.9 BAM [35] 43.4 50.6 47.5 43.4 46.2 49.3 54.2 51.6 49.6 51.2 DCAMA [63] 49.5 52.7 52.8 48.7 50.9 55.4 60.3 59.9 57.5 58.3 HDMNet [58] 43.8 55.3 51.6 49.4 50.0 50.6 61.6 55.7 56.0 56.0 Training-free 32.1 46.6 50.9 48.7 44.6 55.2 49.5 60.7 45.5 52.7
Table 4: Performance comparison of strict few-shot semantic segmentation settings (1 and 5-shot) on COCO-$20^{i}$. We aggregate instance-level predictions to allow comparison with semantic segmentation works. Previous methods results are sourced from [87].
PASCAL VOC Few-Shot Benchmark
The PASCAL-VOC dataset [10] consists of 20 classes. For few-shot evaluation, we adopt the standard approach [16], splitting the classes into three groups, each with 15 base and 5 novel classes. As in prior work [82], we report AP50 results on the novel classes.
Table 2 shows that our method outperforms all previous approaches across all splits, achieving state-of-the-art performance across all splits. This holds for both methods that fine-tune on novel classes and those that do not.
Figure 3: Qualitative results on the COCO val2017 test set under the 10-shot setting (using 10 reference images per class). Bounding box visualisations are thresholded at 0.5. Our method effectively handles multiple overlapping instances in crowded scenes, demonstrating fine-grained semantics and precise localisation. Through semantic-aware soft merging, we avoid duplicate detections and false positives. Best viewed when zoomed in.
4.2 Cross-Domain Few-Shot Object Detection
The CD-FSOD benchmark [14] is designed to evaluate cross-domain few-shot object detection (CD-FSOD) models by addressing challenges in domain shifts and limited data scenarios. It uses COCO [42] as the source training dataset (SD), and six target datasets (TD) — ArTaxOr, Clipart1k, DIOR, DeepFish, NEUDET, and UODD — spanning photorealistic, cartoon, aerial, underwater, and industrial domains with high inter-class variance.
While many approaches fine-tune on a few labeled instances (support set S) from TD before testing on the query set Q, our model is entirely training-free. Thus, we directly evaluate on the six target datasets without any fine-tuning.
Table 3 compares FSOD methods on the CD-FSOD benchmark across 1-shot, 5-shot, and 10-shot settings. Our method sets a new state-of-the-art among training-free approaches and remains competitive with fine-tuned models. These results demonstrate its strong cross-domain generalisation and robustness without requiring retraining.
4.3 Few-shot semantic Segmentation on COCO
Although our method is designed for instance segmentation, we also evaluate it on the COCO-20i Few-Shot Semantic Segmentation benchmark [52]. The 80 COCO classes are divided into four folds [25, 71, 52], each containing 60 base and 20 novel classes. We assess performance on the 20 novel classes under strict 1-shot and 5-shot settings [87]. Results are shown in Table 4.
To adapt our instance segmentation predictions to semantic segmentation, we aggregate all instances of the same class into semantic maps, enabling direct comparison with prior methods. Despite being entirely training-free, our method achieves competitive performance against fine-tuned approaches.
4.4 Variance in reference set
Figure 4: Performance variance across different n-shot settings on the COCO-20i benchmark. We report the mean average precision (mAP) with error bars representing the std over 10 runs with different randomly selected reference images. Results show that variance is higher for lower n-shot settings (e.g., 1–3 shots) due to the greater dependence on reference image selection. As the number of shots increases, the variance decreases, demonstrating the robustness of our training-free method to reference set variations.
Using different reference images leads to variations in results, as performance depends on the quality of the selected reference images for each class.
To quantify this variance, we evaluate our method on the COCO-20i few-shot object detection benchmark using different random seeds to select reference images. Figure 4 shows the standard deviation (std) across 10 runs. We observe that increasing the number of reference images (higher n-shots) reduces result variance, with lower std values. In 1, 2, and 3-shot settings, reference image selection has a more noticeable impact, while for 5 or more shots, the low std demonstrates robustness to reference set variation.
These findings suggest that some reference images are inherently stronger for a given shot setting. Exploring the characteristics of an “optimal” reference set remains an open direction for future work.
5 Conclusions
In this work, we introduced a novel training-free approach for few-shot instance segmentation by integrating SAM’s mask generation capabilities with the fine-grained semantic understanding of DINOv2. Our method uses reference images to construct a memory bank, refines its internal representations with feature aggregation, and performs feature matching for novel instances using cosine similarity and semantic-aware soft merging. We demonstrate that careful engineering of existing frozen foundation models can lead to state-of-the-art performance without the need for additional training: we achieve 36.8% nAP on COCO-FSOD (outperforming fine-tuned methods), 71.2% nAP50 on PASCAL VOC Few-Shot, and strong generalisation across domains (as evidenced on CD-FSOD benchmark). Furthermore, our semantic segmentation results confirm that our approach can be extended beyond instance segmentation by aggregating instance predictions into semantic maps.
For future work, we identify several promising directions: (1) exploring learning-based strategies to automatically find the most informative reference images for 1-5 shot scenarios; (2) addressing DINOv2’s global semantic biases by improving feature localisation, especially for fine-grained tasks; and (3) investigating lightweight finetuning approaches to improve the internal memory bank representations for 1-5 shot scenarios.
6 Acknowledgements
Funding for this research is provided in part by the SENSE - Centre for Satellite Data in Environmental Science CDT studentship (NE/T00939X/1), and an EPSRC New Investigator Award (EP/X020703/1). This work used JASMIN, the UK’s collaborative data analysis environment https://jasmin.ac.uk [36].
Supplementary Material
Figure 5: Cross-domain 5-shot segmentation results using our training-free method. Our approach evaluates diverse datasets across multiple domains, including aerial, underwater, microscopic, and cartoon imagery, without requiring fine-tuning. Results demonstrate the robustness and generalisability of our method.
7 Ablation on aggregation strategies
The improvements of our soft-merging semantic-aware strategy are shown in Tab. 5. Other aggregation variants, such as covariance similarity, instance softmax, score decay, iterative mask refine, attention-guided global average, underperformed.
Aggregation strategy 10-shot nAP Hard-merging (hard threshold of 1 IoS) 31.2 Soft-merging (without semantics) 35.7 Soft-merging (with semantics) 36.6
Table 5: Ablation on aggregation and matching strategies.
8 Ablation on efficiency
Our training-free stages are optimised and lightweight. (1) Memory bank construction (0.1 s/img) is computed once, only requiring to encode $n$ reference images per class with DINOv2. Reference-image features are pre-cached for all following steps. (2) Semantic matching (0.0003 s/img) is a fully parallelised dot product of cosine similarities. (3) Soft-merging (0.006 s/img) uses a parallel implementation of NMS. Tab. 6 shows our method yields significant performance gains compared to Matcher [46] and speeds up SAM default automatic mask generator (AMG) by ×$\times 3$ via efficient point sampling, faster mask filtering, and removal of unnecessary post-processing.
Method Time (sec/img) Matcher [45] 120.014 Training-free (ours) with SAM AMG 3.5092 Training-free (ours) 0.9292
Table 6: Time to process an image on 20 ref. classes with A100.
9 Cross-Domain Few-Shot Object Detection
We provide additional visualisations of the six target datasets in the CD-FSOD benchmark. These datasets span diverse and challenging domains, including photorealistic, cartoon, aerial, underwater, and industrial imagery, each presenting unique distribution shifts. Despite these variations, our method achieves strong performance across all domains without any fine-tuning, demonstrating its remarkable cross-domain generalization. Figure 5 shows an overview of the results for the 6 datasets. Figures 6, 7, 9, 8, 10, 11 display more detailed results for each of the datsets. All results are shown for 5-shot setting, using 5 reference images per category.
Figure 6: 5-shot results on the ArTaxOr dataset.
Figure 7: 5-shot results on the UODD underwater dataset.
Figure 8: 5-shot results on the Fish dataset.
Figure 9: 5-shot results on the Clipart1k dataset.
Figure 10: 5-shot results on the DIOR dataset.
Figure 11: 5-shot results on the NEU-DET dataset.
10 COCO-20i
Figure 12: Visualisation of failure cases of our training-free method on the COCO val2017 set under the 10-shot setting (using 10 reference images per class). Our method sometimes confuses semantically similar classes, such as misclassifying bread as a hot dog or a large armchair as a couch. Additionally, we observe that fine or small objects are occasionally missed, and in highly crowded scenes, our model struggles to detect all instances accurately.
Despite the strong performance of our training-free method across datasets, it also exhibits certain limitations, displayed in Figure 12. A recurring failure mode is the confusion between semantically similar categories, such as bread being misidentified as a hot dog or large armchairs being mistaken for couches. This suggests that our approach could benefit from more fine-grained differentiation or improved selection of reference images. Additionally, detecting small or fine-grained objects remains challenging, as some instances are missed. Finally, in densely crowded scenes, where multiple overlapping objects appear, our model tends to under-detect instances, likely due to occlusions and the complexity of the visual context. These observations highlight areas for future improvement in robust object detection and segmentation under few-shot constraints.